What is the equilibrium constant for a reaction at temperature 78.5 °C if the equilibrium constant at 97.9 °C is 20.64? For this reaction, ΔrH = -7.7 kJ mol-1 .
Answer #1

Given:

T1 = 97.9 oC

=(97.9+273)K

= 370.9 K

T2 = 78.5 oC

=(78.5+273)K

= 351.5 K

K1 = 20.64

Ho = -7.7 KJ/mol

= -7700 J/mol

use:

ln(K2/K1) = (Ho/R)*(1/T1 - 1/T2)

ln(K2/20.64) = (-7700.0/8.314)*(1/370.9 - 1/351.5)

ln(K2/20.64) = -926*(-1.488*10^-4)

K2 = 23.69

Answer: 23.7